(x^2)/(4)+(y^2)/(16)=1
Answer by lwsshak3(11628)

(x^2)/(4)+(y^2)/(16)=1
This is an equation of an ellipse with vertical major axis:
Its standard form: (x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k)=(x,y) coordinates of center
For given equation:
center: ((0,0)
a^2=16
a=√16=4
length of major axis=2a=8
b^2=4
b=2
length of minor axis=2b=4
c^2=a^2-b^2=16-4=12
c=√12 (distance between focal points)